# Reweighted Spectral Partitioning Works: Bounds for Special Graph Classes

### v3

@@ 2.3 Non-Expansive Metric Embeddings
For two metric spaces $(X,d)$ and $(Y,d^{\prime})$ with a metric embedding function $f:X\to Y$, we say that $f$ is non-expansive if $d^{\prime}(f(x),f(y))\leq d(x,y)$ for all $x,y\in X$.
Definition 2.13.
For two metric spaces $(X,d)$ and $(Y,d^{\prime})$ with a metric embedding function $f:X\to Y$, we say the average $p$-distortion of $f$ is
$\frac{\sum_{x,x^{\prime}\in X}d(x,x^{\prime})^{p}}{\sum_{x,x^{\prime}\in X}d^{\prime}(f(x),f(x^{\prime}))^{p}}.$
It’s worth noting here that $\sum_{x,x^{\prime}\in X}\frac{d(x,x^{\prime})^{p}}{d^{\prime}(f(x),f(x^{\prime}))^{p}}\leq\frac{\sum_{x,x^{\prime}\in X}d(x,x^{\prime})^{p}}{\sum_{x,x^{\prime}\in X}d^{\prime}(f(x),f(x^{\prime}))^{p}}$, so an upper bound on the average $p$-distortion also provides an upper bound on this quantity.
This notion of average distortion was studied by Rabinovich [Rab08]. This is a bit different than most typical contexts for metric embeddings, which usually consider “worst-case distortion”. In fact, we can often obtain better bounds for this kind of “average” distortion.
The primary method for obtaining bounds on these average distortion values is by studying yet another quantity:
Definition 2.14.
@@
where $B(x,r)$ denotes the metric ball of radius $r$ centred at $x$.
There are some other definitions of this quantity as well, but this form is suitable for our purposes. In particular, we will not need to use the structure of the modulus of padded decomposability directly – all the relevant results to this work have already been derived.
Theorem 2.15 ([BLR10, Theorem 4.4]).
@@
We are particularly interested in the modulus of padded decomposability for shortest-path metrics on graphs in special classes. We use the following notation:
Definition 2.16.
For a graph $G=(V,E)$, the extremal edge and vertex moduli of padded decomposability on $G$ are defined as the maximum values of $\alpha(V,d_{\omega})$ across all edge-weight functions $\omega:E\to\mathbb{R}_{\geq 0}$ and vertex-weight functions $\omega:V\to\mathbb{R}_{\geq 0}$ (respectively) inducing shortest-path metrics $d_{\omega}$ on $G$. We denote the edge-weighted quantity as $\alpha(G)$, and the vertex-weighted quantity as $\overline{\alpha}(G)$. Note that $\overline{\alpha}(G)\leq\alpha(G)$.
There is a long sequence of results bounding $\alpha(G)$ for various graph classes. A very recent result has subsumed most of them:
Theorem 2.17 ([CF25, Theorem 2]).
Let $G$ be a $K_{h}$-minor-free graph. Then $\alpha(G)\in\operatorname{\mathcal{O}}(\log h)$. Consequently, for any set of edge or vertex weights $\omega$ and an induced shortest-path distance metric $d_{\omega}$, there exists a non-expansive metric embedding function $f:(V(G),d_{\omega})\to\mathbb{R}$ with average 222-distortion at most $\operatorname{\mathcal{O}}((\log h)^{2})$.
Conroy and Filtser outline the recent history of the results subsumed by the above [CF25]. Importantly, the above result also implies that $\alpha(G)=\operatorname{\mathcal{O}}(\log g)$ for a graph of genus $g$.
Lee [Lee17] proved a similar theorem for region intersection graphs that is not immediately subsumed by the above:
Theorem 2.18 ([Lee17]).
For graphs $G$ and $\hat{G}$, where $\hat{G}\in\text{rig}(G)$: If $G$ has no $K_{h}$-minor, then $\overline{\alpha}(G)\in\operatorname{\mathcal{O}}(h^{2})$.
It seems likely that the $\operatorname{\mathcal{O}}(h^{2})$ here could also be reduced to $\operatorname{\mathcal{O}}(\log h)$ as well, but the techniques of both papers are quite involved, so it would take some careful examination. It is also not immediately clear if the $\overline{\alpha}(G)$ could be replaced by $\alpha(G)$.
@@ 3 Bounds via Geometric Intersection Graphs
In this section, we will show that the dimension-restricted dual program $\gamma^{(d)}$ to $\lambda_{2}^{*}$ is intimately related to a certain kind of $d$-dimensional geometric intersection graphs. Using these geometric techniques, we will obtain bounds on both quantities for planar graphs, $k$-nearest neighbour graphs, and bounded-genus graphs. As a consequence, we will recover a new proof of the planar separator theorem. The arguments in this section adapt methods by Spielman and Teng [ST96, ST07], as well as Kelner [Kel04, Kel06b, Kel06a].
@@
$\begin{array}[]{r c c r c l l}\dot{\gamma}^{(d)}(G)&:=&\min_{\begin{subarray}{c}f:V\to\mathbb{R}^{d}\\ s:V\to\mathbb{R}_{\geq 0}\end{subarray}}&\frac{\sum_{v\in V}\left(s(v)\right)^{2}}{\sum_{x\in V}||f(x)||_{2}^{2}}\\ &&\textup{subject to}&\sum_{v\in V}f(v)&=&\hskip 10.00002pt\overline{0}\\ &&&s(u)+s(v)&\geq&\hskip 10.00002pt||f(u)-f(v)||_{2}&\forall uv\in E\end{array}$ (2)
@@
For the first inequality, let $s,f$ be the optimal solution to $\dot{\gamma}^{(d)}(G)$, and take $y(v):=\frac{2s(v)^{2}}{\sum_{x\in V}||f(x)||_{2}^{2}}$ for each $v\in V$. Then, for each edge $uv\in E$,
$y(u)+y(v)=\frac{2s(u)^{2}+2s(v)^{2}}{\sum_{x\in V}||f(x)||_{2}^{2}}\geq\frac{\left(s(u)+s(v)\right)^{2}}{\sum_{x\in V}||f(x)||_{2}^{2}}\geq\frac{||f(u)-f(v)||_{2}^{2}}{\sum_{x\in V}||f(x)||_{2}^{2}},$
so the constraints are satisfied. Moreover $\sum_{v\in V}y(v)=2\cdot\dot{\gamma}^{(d)}(G)$, so we get the inequality $\gamma^{(d)}(G)\leq\sum_{v\in V}y(v)=2\cdot\dot{\gamma}^{(d)}(G)$.
For the second inequality, let $y,f$ be the optimal solution to $\gamma^{(d)}(G)$, and take
$s(v):=\sqrt{y(v)\cdot\sum_{x\in V}||f(x)||_{2}^{2}}$
for each $v\in V$. We perform a similar argument for each $uv\in E$:
$\left[s(u)+s(v)\right]^{2}\geq(s(u))^{2}+(s(v))^{2}=\left[y(u)+y(v)\right]\cdot\sum_{x\in V}||f(x)||_{2}^{2}\geq||f(u)-f(v)||_{2}^{2},$
so the constraints are satisfied after taking the square root of each side. Moreover, $\sum_{v\in V}s(v)^{2}=\gamma^{(d)}(G)$, so we get the inequality $\dot{\gamma}^{(d)}(G)\leq\sum_{v\in V}s(v)^{2}=\gamma^{(d)}(G)$. ∎
Importantly, $\dot{\gamma}^{(d)}(G)$ gives us a new geometric interpretation of the problem: A solution satisfying the constraints corresponds to a representation of $G$ as a subgraph of a geometric intersection graph: For each vertex $v$, create a $d$-dimensional ball $B_{v}$ of radius $s(v)$ centered at $f(v)$. Construct a new graph $H=(V,F)$ so that $uv\in F$ if and only if $B_{u}\cap B_{v}\neq\emptyset$. Then $H\supset G$. As a consequence, if we can bound the sum of squared radii for such a geometric representation, with the normalization constraints $\sum_{x\in V}||f(x)||_{2}^{2}=1$ and $\sum_{x\in V}f(x)=\overline{0}$, then we can bound $\gamma^{(d)}(G)$.
One particularly important natural case of a graph class that admits such representations are planar graphs, in which such a representation corresponds to special form of the well-celebrated circle packing theorem. We start with a definition:
Definition 3.2.
A planar circle packing $\mathcal{P}$ of a simple undirected graph $G$ is a set of (possibly overlapping) circles $\{C\}_{v\in V(G)}$ in the complex plane $\mathbb{C}$ such that $uv\in E(G)$ if and only if the circles $C_{u}$ and $C_{v}$ are mutually tangent (with disjoint interiors). If all pairs of circles also have pairwise disjoint interiors, then we say that $\mathcal{P}$ is univalent.
Figure 1: Two examples planar circle packings. The left planar circle packing is univalent, while the right is not.
See Figure 1 for examples of univalent and non-univalent circle packings. We can now state the theorem:
@@
Let $G$ be a simple undirected graph. Then $G$ has a univalent planar circle packing if and only if $G$ is a planar graph. Moreover, if $G$ is a maximal planar graph, then this circle packing is unique up to Möbius transformations.
Spielman and Teng [ST96, ST07] made use of this theorem to bound $\lambda_{2}(G)$ for planar graphs $G$, hence giving a proof of a weaker version of the planar separator theorem via Theorem 1.1. We will use a similar argument to bound $\lambda_{2}^{*}(G)$, allowing us to give a new proof of the full planar separator theorem via Theorem 1.4. We will make use of one of their theorems to accomplish this:
@@
[…]
This theorem statement is slightly weaker than the statement used by Spielman and Teng, who also described the structure of the homeomorphism as a stereographic projection. However, it is sufficient for our purposes. In particular, it gives us a method of “normalizing” a geometric intersection graph of balls in $\mathbb{R}^{d}$ to be a geometric intersection graph of (geodesic) balls in $S^{d}$. Importantly, since the theorem gives a homeomorphism, it preserves the ply of each individual point, which is the total number of balls containing that point. From this, we can now obtain the desired bound on $\lambda_{2}^{*}$ for planar graphs:
Theorem 3.5.
Let $G$ be a planar graph with $n$ vertices and maximum degree $\Delta$. Then $\frac{\lambda_{2}(G)}{\Delta}\leq\lambda_{2}^{*}(G)\leq\gamma^{(3)}(G)\leq\frac{8}{n}$. Hence, $\psi(G)^{2}\lesssim\frac{1}{n}$.
Note that this theorem is a strengthening of both Theorem 2.1 and the planar separator theorem. The proof method will also be analogous to that of Spielman and Teng [ST96, ST07, proof of Theorem 3.3].
@@
By Theorem 3.3 and Theorem 3.4, there exists a representation of $G$ as a circle packing on the sphere $S^{2}\subset\mathbb{R}^{3}$. Let the centers be given by $f:V\to S^{2}\subset\mathbb{R}^{3}$ (i.e., $||f(v)||_{2}^{2}=1$ for each $v\in V$) and let the radii be given by $s:V\to\mathbb{R}_{\geq 0}$. We claim first that $f,s$ form a feasible solution to $\dot{\gamma}^{(3)}(G)$: The constraint $\sum_{v\in V}f(v)=\overline{0}$ is exactly the statement that the centroid of the centres is the origin, which is given by Theorem 3.4. The other constraints follow from the facts that two balls intersect if and only if they share an edge, and that $||f(u)-f(v)||_{2}$ is bounded above by the geodesic distance along $S^{2}$ between $f(u)$ and $f(v)$. Next, we bound the objective value: It follows from the statement of Theorem 3.3 that the ply of all points in $S^{2}$ except for a set of measure 0 is at most 111. Hence, the total area of all balls is bounded by the area of $S^{2}$ itself, which is known to be $4\pi$. The area of the ball on $S^{2}$ for vertex $v$ is bounded above by $\pi(s(v))^{2}$, since $s(v)$ is the (geodesic) radius of the ball, so $\sum_{v\in V}\pi(s(v))^{2}\leq 4\pi$. Therefore, we get a bound on the objective value of
$\frac{\sum_{v\in V}s(v)^{2}}{\sum_{x\in V}||f(x)||_{2}^{2}}\leq\frac{4}{n}.$
@@
3.1 Geometric Intersection Graphs and Nearest-Neighbour Graphs
Spielman and Teng [ST96, ST07] also used higher-dimensional geometric intersection graphs with similar properties to bound $\lambda_{2}$ for a few other graph classes. In particular, they obtained bounds for $d$-dimensional $k$-ply neighbourhood systems, and $d$-dimensional $k$-nearest-neighbour graphs.
We can strengthen another key result of theirs:
Theorem 3.6.
[…]
$\frac{\lambda_{2}(G)}{\Delta}\leq\lambda_{2}^{*}(G)\leq\gamma^{(d+1)}(G)\leq 2\cdot\left(\frac{A_{d}}{V_{d}}\cdot\frac{k}{n}\right)^{\frac{2}{d}}<2\cdot\left(2\pi\cdot\left(d+2\right)\right)^{\frac{1}{d}}\cdot\left(\frac{k}{n}\right)^{\frac{2}{d}}\lesssim\left(\frac{k}{n}\right)^{\frac{2}{d}}.$
Consequently, $\psi(G)\lesssim\sqrt{\min\{d,\log\Delta\}}\cdot\left(\frac{k}{n}\right)^{\frac{1}{d}}$. Moreover, reweighted spectral partitioning will find, in polynomial time, a sparse vertex cut $S$ with $\frac{|\partial S|}{|S|}\lesssim\sqrt{\log\Delta}\cdot\left(\frac{k}{n}\right)^{\frac{1}{d}}$.
Note that this theorem is essentially a further generalization of Theorem 3.5.

### v4

@@ 4 A Refined Cheeger Inequality
For a graph $G$, we define the $d$-dimensional $L^{p}$-extremal embedded spread as:
@@
For the purposes of this section, $Q^{(1)}_{2}$ will be a bit easier to work with than $\gamma^{(1)}$, and the more general definition will also be useful in a later section when we try to relate $\gamma^{(1)}$ and the $L^{2}$-extremal spread $\overline{s}_{2}$. $Q^{(d)}_{2}(G)$ and $\gamma^{(d)}(G)$ are related as follows:
Lemma 4.4.
For a graph $G$, $\gamma^{(d)}(G)=\frac{2n}{\overline{Q}^{(d)}_{2}(G)}$
Proof.
We claim that for any function $f:V\to\mathbb{R}^{d}$ for which $\sum_{v\in V}f(v)=\overline{0}$, it holds that
@@
To prove this claim, let $f_{i}(v)$ denote the $i$th coordinate of the vector $f(v)$. Then,
[…]
$=\sum_{u,v\in V}\sum_{i\in[d]}\left[f_{i}(u)^{2}+f_{i}(v)^{2}-2f_{i}(u)f_{i}(v)\right]$
$=\sum_{i\in[d]}\sum_{u\in V}\sum_{v\in V}\left[f_{i}(u)^{2}+f_{i}(v)^{2}-2f_{i}(u)f_{i}(v)\right]$
$=\sum_{i\in[d]}\sum_{u\in V}\left[\sum_{v\in V}\left(f_{i}(u)^{2}+f_{i}(v)^{2}\right)-2f_{i}(u)\sum_{v\in V}f_{i}(v)\right]$
$=\sum_{i\in[d]}\sum_{u\in V}\sum_{v\in V}\left[f_{i}(u)^{2}+f_{i}(v)^{2}\right]$
$=\sum_{i\in[d]}2n\sum_{x\in V}f_{i}(x)^{2}$
[…]
Therefore,
[…]
No vector f(u)f(v)f(u)-f(v) is affected by translation, so the constraint $\sum_{v\in V}f(v)=\overline{0}$ is unnecessary, and thus the reciprocal of the inner fractional program here is exactly $\overline{Q}_{2}^{(d)}$. ∎
The key lemma is as follows:
Lemma 4.5.
For any graph $G$, and any $d,p\geq 1$, $\overline{Q}_{p}^{(d)}(G)\leq C_{p}\cdot\left[\alpha(G)\right]^{p}\cdot\overline{Q}_{p}^{(1)}(G),$ where $C_{p}$ is a constant dependent only on $p$. Moreover, given a feasible solution to $\overline{Q}_{p}^{(d)}(G)$ and a padded partition scheme with padding $\alpha$, a corresponding feasible solution to $Q_{p}^{(1)}(G)$ satisfying this inequality can be found using Algorithm 2 with high probability.
Proof.
Let $f,y$ be the optimal solution to $Q_{p}^{(d)}$. Let $\omega(v):=y(v)^{\frac{1}{p}}$. Let $d_{\omega}$ be the shortest-path metric over $G$ that uses $\omega$ as vertex weights, so a path $v_{1},\dots,v_{k}$ has weight $\frac{1}{2}\omega(v_{1})+\omega(v_{2})+\cdots+\omega(v_{k-1})+\frac{1}{2}\omega(v_{k})$. Biswal, Lee, and Rao [BLR10, Theorem 4.4] showed that applying Algorithm 2 (or a generalization of this algorithm for arbitrary $p$) embeds the metric $(V,d_{\omega})$ into $\mathbb{R}$ with a function $f^{\prime}:V\to\mathbb{R}$ so that:
  • For every pair of vertices $u$ and $v$, $|f^{\prime}(u)-f^{\prime}(v)|\leq d_{\omega}(u,v)$.
  • With high probability, $C_{p}^{\prime}\cdot\alpha(V,d_{\omega})^{p}\sum_{u,v\in V}|f^{\prime}(u)-f^{\prime}(v)|^{p}\geq\sum_{u,v\in V}d_{\omega}(u,v)^{p}$, for some positive constant $C_{p}^{\prime}$ depending only on $p$.
To verify that $f^{\prime},y$ forms a feasible solution to $Q_{p}^{(1)}$, consider an edge $uv\in E$. For such an edge, the shortest path is always to use the edge, so $d_{\omega}(u,v)=\frac{1}{2}\left[\omega(u)+\omega(v)\right]$. Therefore, $|f^{\prime}(u)-f^{\prime}(v)|^{p}\leq d_{\omega}(u,v)^{p}\leq\left(\frac{\omega(u)+\omega(v)}{2}\right)^{p}\leq\frac{\omega(u)^{p}+\omega(v)^{p}}{2}<y(u)+y(v)$, where the penultimate inequality follows from Jensen on the convex mapping $x\mapsto x^{p}$.
We now bound the objective value. For each $u,v\in V$, there is some path $u=w_{1},w_{2},\dots,w_{k}=v\in V$ with
$d_{\omega}(u,v)$ $=\frac{1}{2}\omega(w_{1})+\omega(w_{2})+\cdots+\frac{1}{2}\omega(w_{k})$
$=\frac{1}{2}\left[\left(\omega(w_{1})+\omega(w_{2})\right)+\left(\omega(w_{2})+\omega(w_{3})\right)+\cdots+\left(\omega(w_{k-1})+\omega(w_{k})\right)\right].$
[…]
Picking $C_{p}:=2C_{p}^{\prime}$ then gives the stated inequality. ∎
The combination of Lemma 4.4 and Lemma 4.5 therefore completes the proof of Theorem 1.4.
@@ 5 Bounds via Geometric Intersection Graphs
In this section, we prove Propositions 5.5, 2.5, 2.6 and 2.7. Each of these is a bound on $\gamma^{(d)}$ for a different graph class (and a few different values of $d$). In particular, each of these bounds uses the relationships these classes have with a few special kinds of $d$-dimensional ball-intersection graphs, and how such ball-intersection graphs relate directly to $\gamma^{(d)}$.
@@
[…]
@@
[…]
The special conditions we need for our ball-intersection graphs all involve the concept of ply: For a set of balls $\mathcal{B}$ in $\mathbb{R}^{d}$, a point $p\in\mathbb{R}^{d}$ is said to have ply $k$ if there are exactly $k$ balls in $\mathcal{B}$ that contain $p$. Then, the ply of $\mathcal{B}$ itself is said to be the maximum ply value of any point not contained in the boundary of some ball. Note that, due to this boundary condition, sets of balls with even ply $1$ can still contain intersections. For planar graphs and the geometric graph classes, we will be able to work with bounded ply for the full set of balls. However, for bounded-genus graphs, we instead will use set of balls containing a very large subset that has bounded ply. See Figure 2 for some examples of balls with bounded ply, and their corresponding geometric intersection graphs. A set of $2$-dimensional balls with ply $1$ is also called a (univalent) circle packing.
Figure 2: Two examples sets of balls, and the corresponding geometric intersection graphs. The left example has ply $1$, while the right example has ply $2$.
It will also be helpful for the results in this section to note the following:
Lemma 5.2.
[OTZ22, Proof of Proposition 2.9]. For a graph $G$ and any value $d\geq 1$, $\gamma^{(1)}(G)\lesssim d\gamma^{(d)}(G)$. Moreover, given a feasible solution to $\gamma^{(d)}(G)$, a feasible solution to $\gamma^{(1)}(G)$ satisfying this inequality can be found in polynomial time.
5.1 Planar Graphs
Planar graphs admit a very well-structured representation as a ball-intersection graph. This comes from the celebrated circle packing theorem:
@@ 5.1 Planar Graphs
Let $G$ be a simple undirected graph. Then $G$ is the geometric intersection graph of $2$-dimensional balls with ply $1$ (a circle packing) if and only if $G$ is a planar graph. Moreover, if $G$ is a maximal planar graph, then this representation is unique up to Möbius transformations.
Spielman and Teng [ST96, ST07] made use of this theorem to bound $\lambda_{2}(G)$ for planar graphs $G$, hence giving a “spectral” proof of a weaker version of the planar separator theorem via Theorem 2.1. We will use a similar argument to bound $\gamma^{(1)}(G)$, hence also implying a “spectral”-type of proof of the planar separator theorem in full generality via Theorem 1.4. We will make use of one key result of Spielman and Teng to accomplish this:
@@
Let $B_{1},\dots,B_{n}$ be a collection of balls in $\mathbb{R}^{d}$ with centres $c_{1},\dots,c_{n}$, so that no point $x\in\mathbb{R}^{d}$ is contained in $\left\lceil\frac{n}{2}\right\rceil$ of the balls. Then there is a homeomorphism $\alpha$ from $\mathbb{R}^{d}$ to a subset of the sphere $S^{d}$, [^4] We use the standard topological convention of $S^{d}$ as the $d$-dimensional surface of the $(d+1)$-dimensional unit ball embeddable in $\mathbb{R}^{d+1}$. This differs from the notation of Spielman and Teng, who used $S^{d}$ to denote the (d1)(d-1)-dimensional surface of the $d$-dimensional unit ball. so that $\alpha(B_{i})$ is exactly a geodesic ball (or sphere cap) in $S^{d}$ with center $\alpha(c_{i})$, and moreover so that the centroid of the values $\alpha(c_{1}),\dots,\alpha(c_{n})$, in the natural representation of $S^{d}$ as surface of the unit ball of $d+1$ dimensions, is exactly the origin.
This theorem statement is slightly weaker than the statement used by Spielman and Teng, who also described the structure of the homeomorphism as a stereographic projection. However, this statement is sufficient for our purposes. In particular, it gives us a method of normalizing a geometric intersection graph of balls in $\mathbb{R}^{d}$ to be a geometric intersection graph of (geodesic) balls in the sphere $S^{d}$. Importantly, since the theorem gives a homeomorphism, it preserves the ply of each individual point. From this, we can now obtain the desired bound on $\gamma^{(1)}$ for planar graphs:
Proposition 5.5 (Extended restatement of Proposition 2.4).
Let $G$ be a planar graph with $n$ vertices and maximum degree $\Delta$. Then,
$\gamma^{(1)}(G)\lesssim\gamma^{(3)}(G)\leq\frac{8}{n}.$
Hence, $\psi(G)^{2}\lesssim\frac{1}{n}$.
The method to prove this will be analogous to a proof of Spielman and Teng [ST96, ST07, proof of Theorem 3.3].
@@
By Theorem 5.3 and Theorem 5.4, there exists a representation of $G$ as a circle packing on the sphere $S^{2}\subset\mathbb{R}^{3}$. Let the centers be given by $f:V\to S^{2}\subset\mathbb{R}^{3}$ (i.e., $f(v)||_{2}^{2}=1$ for each $v\in V$). For each $v\in V$, let $s(v)$ be the longest Euclidean distance from $f(v)$ to a point in its geodesic ball on $S^{2}$, embedded in $\mathbb{R}^{3}$. See Figure 3 for an example of these values. Note that the surface area of the geodesic ball is given exactly by $\pi s(v)^{2}$. We claim first that $f,s$ form a feasible solution to $\dot{\gamma}^{(3)}(G)$: The constraint $\sum_{v\in V}f(v)=\overline{0}$ is exactly the statement that the centroid of the centres is the origin, which is given by Theorem 5.4. The other constraints follow from the fact that two balls intersect if and only if they share an edge. Next, we bound the objective value: It follows from the statement of Theorem 5.3 that the ply of all points in $S^{2}$ except for a set of measure 0 is at most $1$. Hence, the total area of all balls is bounded by the area of $S^{2}$ itself, which is $4\pi$. That is, $\sum_{v\in V}\pi s(v)^{2}\leq 4\pi$. Therefore, we get a bound on the objective value of
[…]
@@
$f(v)$$s(v)$
Figure 3: Illustration of a geodesic ball (sphere cap) on the unit sphere $S^{2}$ embedded in $\mathbb{R}^{3}$, along with its corresponding centre $f(v)$. The distance $s(v)$ is the longest Euclidean distance from $f(v)$ to any point in the ball.
5.2 Ball-Intersection Graphs and Nearest-Neighbour Graphs
Spielman and Teng [ST96, ST07] also considered higher-dimensional geometric graphs with similar properties to planar graphs, and bounded $\lambda_{2}$ for these classes. In particular, they obtained bounds for geometric intersection graphs of $d$-dimensional balls with ply $k$, and $d$-dimensional $k$-nearest-neighbour graphs. We can use similar techniques to prove Propositions 2.5 and 2.6.
Unfortunately, their proof also contains a minor error. Specifically, they used an inequality that does not hold for $d\geq 3$ dimensions: Let $A_{d}$ denote the area of the unit $d$-sphere $S^{d}$, and let $V_{d}$ denote the volume of the unit $d$-ball $B^{d}$ (both of these are $d$-dimensional measures). [^5] We again use the topological convention of the $d$-sphere $S^{d}$ as the $d$-dimensional surface of the $(d+1)$-dimensional unit ball. Let $C$ be a sphere cap (geodesic ball) on the unit $d$-sphere $S^{d}$ embedded in $\mathbb{R}^{d+1}$. Let $f$ be the centre of $C$ on $S^{d}$. Let $r$ be the maximum Euclidean distance from $f$ to a point in $C$. They claimed that $V_{d}r^{d}\leq\text{volume}(B)$. This is true for $d\leq 2$. In fact, we implicitly make use of this fact in the previous subsection. However, we will now provide a counter-example for $d=3$:
Example 5.6.
For $d=3$ dimensions, any sphere cap $C$ that covers strictly less than half the hypersurface of $S^{3}$ has $3$-dimensional surface area strictly less than the volume of the $3$-dimensional ball $V_{3}$ of radius $r$, where $r$ is the maximum Euclidean distance through $\mathbb{R}^{4}$ from the centre of the sphere cap to any other point inside it.